OpenStudy (anonymous):

A centrifuge is used to test space pilots. The centrifuge spins with a centripetal acceleration of 3.04 g. If the length of the centrifuge's arm is 21 m, what is the speed of the centrifuge? A. 21 m/s B. 25 m/s C. 27 m/s D. 29 m/s

OpenStudy (anonymous):

here the force responsible for making the centrifuge spin about a point is the centripetal force.according to newton's second law, F=mxa but centripetal force is given by F=(mv^2)/r equate these two equations and solve for v r=radius of the circle in which the object is moving

OpenStudy (anonymous):

a = 3.04 g = v^2 / r g = 9.8, m/s^2 solve for a

OpenStudy (anonymous):

I mean, solve for v.

OpenStudy (anonymous):

what's the equation id use to solve for v? @douglaswinslowcooper

OpenStudy (anonymous):

v = sqrt(3.04 r g) You should be able to do this kind of solution now. Put the unknown on one side of the equation, then do what is needed to get it as , for example, v = .....

OpenStudy (anonymous):

I got 29 m/s (d) Is that correct? @douglaswinslowcooper

OpenStudy (anonymous):

I got 25 = sqrt[(3.04)(9.8)(21)]